# Cyclotomic generating functions

### v2

@@ 5. Analytic considerations
(5.1) $\phi_{\mathcal{X}}(z)=f(e^{iz})/f(1)=\Phi_{i_{1}}(e^{iz})\Phi_{i_{2}}(e^{iz})\cdots\Phi_{i_{k}}(e^{iz})/f(1)=\prod_{j=1}^{k}\frac{[a_{j}]_{e^{iz}}}{[b_{j}]_{e^{iz}}}$
@@
(5.2) $\phi_{\mathcal{X}^{*}}(z)$ $=\mathbb{E}[e^{iz\mathcal{X}^{*}}]=\mathbb{E}[e^{iz(X-\mu)/\sigma}]=e^{-iz\mu/\sigma}\mathbb{E}[e^{izX/\sigma}]$
@@
(5.3) $\log\phi_{\mathcal{X}^{*}}(z)$ $=\sum_{d=1}^{\infty}(-1)^{d}\kappa_{2d}^{*}(f)\frac{z^{2d}}{(2d)!}$
@@
since $\kappa_{2d}^{*}(f)=\kappa_{2d}(f)/\kappa_{2}(f)^{d}$ and $\kappa_{d}(f)$ is given by (2.5) for all positive integers $d$. Furthermore, $(-1)^{d}\kappa_{2d}^{*}(f)$ is a non-positive real number by 2.9.
5.1. CGF characteristic functions
Consider the set of standardized characteristic functions of random variables associated to cyclotomic generating functions. As mentioned above, it suffices to consider only basic CGF’s, so consider the set of all standardized CGF characteristic functions on the complex plane,
@@ 5.1. CGF characteristic functions
Given $\phi_{\mathcal{X}^{*}}(z)\in\mathcal{C}_{\operatorname{CGF}}$, let $f(q)\in\Phi^{+}$ be the associated CGF, and let $\mathcal{X}$ be the corresponding random variable with $\mathbb{E}[q^{\mathcal{X}}]=f(q)/f(1)$, mean $\mu$ and standard deviation $\sigma$. Let $\tilde{f}(q)=q^{-\mu}f(q)/f(1)\in\mathbb{R}_{\geq 0}[q^{\pm 1/2}]$. Then, by (5.2)
(5.4) $\phi_{\mathcal{X}^{*}}(z)=e^{-iz\mu/\sigma}f(e^{iz/\sigma})/f(1)=\tilde{f}(e^{iz/\sigma}).$
@@
From Montel’s theorem in complex analysis, $\mathcal{C}_{\operatorname{CGF}}$ is a normal family of entire functions if and only if it is bounded on all complex disks $|z|\leq R$. Hence we will bound $|\phi_{\mathcal{X}^{*}}(z)|$ in terms of $R$. Note that $f(q)=1$ if and only if $\sigma=0$ by 3.9(iii), in which case $\phi_{\mathcal{X}^{*}}(z)=1$ is bounded by any function greater than 1. So, we will assume $f(q)$ is not constant and $\sigma>0$. By [HZ15, Lem. 2.8], for all real $t\geq 0$,111[HZ15, Lem. 2.8] is incorrectly stated for all $t\in\mathbb{R}$, though one of the last steps in the argument requires $t\geq 0$.
(5.5) $\mathbb{E}[e^{t\mathcal{X}^{*}}]\leq\exp\left(\frac{3}{2}t^{2}e^{2t/\sigma}\right).$
Since $\mathbb{E}[e^{z\mathcal{X}^{*}}]=\phi_{\mathcal{X}^{*}}(-iz)$, we can use this inequality to bound $|\phi_{\mathcal{X}^{*}}(z)|$ for all complex $|z|\leq R$ as follows.
For all $|z|\leq R$, we claim that $|\tilde{f}(e^{iz/\sigma})|\leq 2\tilde{f}(e^{R/\sigma})$. Indeed, since $f(q)\in\Phi^{+}$, we have $\tilde{f}(q)=\sum_{k=-N}^{N}a_{k}q^{k/2}$ for $a_{k}\geq 0$ satisfying $a_{k}=a_{-k}$. Therefore,
$\left|\tilde{f}(e^{iz/\sigma})\right|$ $=\left|\sum_{k}a_{k}e^{izk/2\sigma}\right|\leq\sum_{k}a_{k}\left|e^{izk/2\sigma}\right|=\sum_{k}a_{k}e^{\operatorname{Re}izk/2\sigma}$
$\leq\sum_{k}a_{k}e^{R|k|/2\sigma}\leq\sum_{k}a_{k}(e^{Rk/2\sigma}+e^{-Rk/2\sigma})=\sum_{k}a_{k}e^{Rk/2\sigma}+\sum_{k}a_{-k}e^{-Rk/2\sigma}$
@@
By 3.9(iii), $\sigma^{2}\geq 1/4$ is bounded away from 00 since we assumed $f(q)$ is not constant. Therefore, by the claim, (5.4), and (5.5), we have the required uniform bound
$|\phi_{\mathcal{X}^{*}}(z)|=|\tilde{f}(e^{iz/\sigma})|\leq 2\tilde{f}(e^{R/\sigma})=2\phi_{\mathcal{X}^{*}}(-iR)=2\mathbb{E}[e^{R\mathcal{X}^{*}}]\leq 2\exp\left(\frac{3}{2}R^{2}e^{8R}\right).$
Theorem 5.2 (Converse of Frechét–Shohat for CGF’s).
Suppose $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ is a sequence of random variables corresponding to cyclotomic generating functions such that $\mathcal{X}_{n}^{*}\Rightarrow\mathcal{X}$ for some random variable $\mathcal{X}$. Then $\mathcal{X}$ is determined by its moments and, for all $d\in\mathbb{Z}_{\geq 1}$,
@@
By 5.1, we may pass to a subsequence and assume $\phi_{\mathcal{X}_{n}^{*}}(z)$ converges uniformly on compact subsets of $\mathbb{C}$. Hence, they converge to an entire function $\phi(z)$. On the other hand, by Lévy continuity, $\phi_{\mathcal{X}_{n}^{*}}(t)\to\phi_{\mathcal{X}}(t)$ pointwise for all $t\in\mathbb{R}$. Thus, $\phi_{\mathcal{X}}(t)=\phi(t)$ for $t\in\mathbb{R}$.
@@
We may now prove 1.8 from the introduction.
Proof of 1.8.
Suppose $\mathcal{X}_{N}^{*}\Rightarrow\mathcal{X}$ converges in distribution. Then the result follows by 5.2. Conversely, suppose $\lim_{N\to\infty}\mu_{d}^{\mathcal{X}_{N}^{*}}$ exists and is finite for all $d\in\mathbb{Z}_{\geq 1}$. By 5.1, we may pass to a subsequence on which $\phi_{\mathcal{X}_{N}^{*}}(z)$ converges uniformly on compact subsets of $\mathbb{C}$. The limiting entire function $\phi(z)$ then has power series coefficients determined by the $\mu_{d}:=\lim_{N\to\infty}\mu_{d}^{\mathcal{X}_{N}^{*}}$, so the limit is independent of the subsequence we passed to and $\lim_{N\to\infty}\phi_{\mathcal{X}_{N}^{*}}(t)=\phi(t)$ for all $t\in\mathbb{R}$. Moreover, $\phi(t)$ is continuous at 00, so there exists $\mathcal{X}$ such that $\mathcal{X}_{N}^{*}\Rightarrow\mathcal{X}$ by [Bil95, Cor. 26.1]. ∎
5.2. Formal cumulants and cumulant generating functions
The formal cumulants of cyclotomic polynomials have the following explicit growth rate.
Lemma 5.4.
For each fixed $d\geq 1$, the formal cumulants of the cyclotomic polynomials for $n>1$ satisfy
(5.6) $\frac{(2d)!}{d}\left(\frac{n}{2\pi}\right)^{2d}\leq|\kappa_{2d}|\leq\frac{|B_{2d}|n^{2d}}{2d}.$
@@ 5.2. Formal cumulants and cumulant generating functions
Applying the expression for the zeta function at positive even integers gives
@@
$\geq\frac{|B_{2d}|}{2d}\frac{1}{\zeta(2d)}$
@@
While cyclotomic generating functions are typically given in rational form, 5.4 allows their cumulants to be described in terms of the cyclotomic form as follows. One may for instance use 3.5 to describe the $q$-hook formula in cyclotomic form combinatorially.
Corollary 5.5.
If $f(q)=\prod_{j=1}^{k}\Phi_{n_{j}}(q)$, then for each $d\geq 1$ we have
@@
The asymptotic behavior of sequences of CGF random variables can be determined by the standardized cumulant generating function. For instance, $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ is asymptotically normal if and only if the corresponding sequence of standardized cumulants $(\kappa_{d}^{(N)})^{*}\to 0$ for all $d\geq 3$ as $N\to\infty,$ in which case by (5.3), $\log\phi_{\mathcal{X}_{N}^{*}}(t)\to-t^{2}/2$ for all $t\in\mathbb{R}$ as $N\to\infty$. Standardized second characteristic functions of CGF’s are particularly “well-behaved” in the following sense.
Proposition 5.6.
@@
(5.7) $\log\phi_{\mathcal{X}^{*}}(t)\leq-\frac{t^{2}}{2}=\log\phi_{\mathcal{N}(0,1)}(t).$
@@
By 5.1, $\phi_{\mathcal{X}}(z)$ is entire, so the singularities of $\log\phi_{\mathcal{X}}(z)$ come about from the zeros of $\phi_{\mathcal{X}}(z)$. To determine the zeros, it suffices to assume $\mathcal{X}$ is determined by $f(q)=\frac{\prod_{k=1}^{m}[a_{k}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}\in\Phi^{+}$ since the results pertain to the standardized random variable corresponding to $\mathcal{X}$. Furthermore, since $\phi_{\mathcal{X}}(z)=1$ if $f(q)=1$, which satisfies both claims, we may assume that $m=\max\{a_{k}\}>1$ and $m$ does not appear in the denominator multiset. Hence, $\Phi_{m}(q)$ is a factor of $f(q)$. Furthermore, from the product formula (2.14), we observe that the closest zero of $\phi_{\mathcal{X}}(z)=f(e^{iz})$ to the origin occurs at $z=2\pi/m$. Hence $\phi_{\mathcal{X}}(z)$ is analytic on the simply connected domain $|z|<2\pi/m$, so $\log\phi_{\mathcal{X}}(z)$ is analytic on the disk $|z|<2\pi/m$.
@@
Since $\Phi_{m}(q)$ is a factor of $f(q)$, we have $0<\kappa_{2}(\Phi_{m})\leq\sigma^{2}$ by 5.4 and 5.5. By the lower bound in (5.6), $\sqrt{\kappa_{2}(\Phi_{m})}\geq\frac{m}{\pi\sqrt{2}}$. Hence if $|z|<\sqrt{2}$,
@@
(5.8) $\log\phi_{\mathcal{X}^{*}}(z)=-\sum_{d\geq 1}\frac{(-1)^{d-1}\kappa_{2d}}{(2d)!}\left(\frac{z}{\sigma}\right)^{2d}$
to compute $\log\phi_{\mathcal{X}^{*}}(t)$ for real $-\sqrt{2}<t<\sqrt{2}$. The inequality (5.7) comes from truncating this power series expansion after the first term and applying 2.9 to see that all terms on the right side are nonpositive. ∎
6. CGF monoids and related open problems
As mentioned in Definition 1.3, the set of basic cyclotomic generating functions forms a monoid under multiplication, $\Phi^{+}$. We also consider the larger monoid generated by all cyclotomic polynomials. Since many families of polynomials of interest are also either unimodal or log-concave with no internal zeros, we consider these submonoids as well, together with another variant using the Gale order on multisets. We conclude with a monoid associated to Hilbert series of polynomial rings quotiented by regular sequences and several open problems.
6.1. Basic, unimodal, and log-concave CGF monoids
Definition 6.1.
@@ 6.1. Basic, unimodal, and log-concave CGF monoids
Lemma 6.2.
We have $\Phi^{\operatorname{lcc}}\subset\Phi^{\operatorname{uni}}\subset\Phi^{+}\subset\Phi^{\pm}$, and for each monoid $\mathcal{M}\in\{\Phi^{\operatorname{lcc}},\Phi^{\operatorname{uni}},\Phi^{+},\Phi^{\pm}\}$ the following facts hold.
@@
For (iv), first note that $f(1)\geq 1\neq 0$ for $f\in\Phi^{+}$, so $\Phi_{1}(q)=q-1$ is not a factor of $f$. Further recall that $\Phi_{n}(0)=\Phi_{n}(1)=1$ for $n\geq 2$ not a prime power. Hence, if $f$ had no factors of the form $\Phi_{p^{k}}(q)$, we would have $f(0)=f(1)=1$, forcing $f(q)=1$ since the coefficients of $f$ are non-negative. The forwards implication in (v) is given by (iv), and the backwards implication follows from the fact that
@@
Finally, for (vi), $\varphi(n)$ is even for all $n\geq 3$, so any product of cyclotomic polynomials of odd degree must contain $\Phi_{2}$. ∎
Example 6.3.
@@
Intuitively, multiplying by $\Phi_{p}(q)=1+q+\cdots+q^{p-1}$ for $p$ prime tends to smooth out chaotic coefficients. Computationally, it appears to be necessary to include such a factor in unimodal or log-concave CGF’s. More precisely, we conjecture the following analogue of 6.2(iv), which has been checked up to degree $50$.
Conjecture 6.4.
@@
We use the cyclotomic form for basic CGF’s and the fact that we know a lower bound for the degree of the cyclotomic polynomial $\Phi_{n}(q)$ to find all basic CGF’s of a given degree. Some initial data and OEIS identifiers [OEI23] can be found in Section 7. The sequence corresponding to $\Phi_{n}^{\pm}$ [OEI23, A120963] was created in 2006. We added the corresponding pages for the other cyclotomic monoids, so we believe the other cyclotomic monoids have not been studied in the literature.
6.2. Gale order and the associated CGF monoid
Given two multisets of the same size, say $A=\{a_{1}\leq a_{2}\leq\cdots\leq a_{m}\}$ and $B=\{b_{1}\leq b_{2}\leq\cdots\leq b_{m}\}$ both sorted into increasing order, we say $A\preceq B$ in Gale order provided $a_{k}\leq b_{k}$ for all $1\leq k\leq m$. This partial order is known by many other names; we are following [ARW16] for consistency. Gale studied this partial order in the context of matroids on $m$-subsets of $\{1,2,\dots,n\}$ in the 1960’s [Gal68].
Definition 6.7.

### v3

@@ 5 Analytic considerations
$\phi_{\mathcal{X}}(z)=f(e^{iz})/f(1)=\Phi_{i_{1}}(e^{iz})\Phi_{i_{2}}(e^{iz})\cdots\Phi_{i_{k}}(e^{iz})/f(1)=\prod_{j=1}^{k}\frac{[a_{j}]_{e^{iz}}}{[b_{j}]_{e^{iz}}}$ (5.1)
@@
$\phi_{\mathcal{X}^{*}}(z)$ $=\mathbb{E}[e^{iz\mathcal{X}^{*}}]=\mathbb{E}[e^{iz(X-\mu)/\sigma}]=e^{-iz\mu/\sigma}\mathbb{E}[e^{izX/\sigma}]$ (5.2)
@@
$\log\phi_{\mathcal{X}^{*}}(z)$ $=\sum_{d=1}^{\infty}(-1)^{d}\kappa_{2d}^{*}(f)\frac{z^{2d}}{(2d)!}$ (5.3)
@@
since $\kappa_{2d}^{*}(f)=\kappa_{2d}(f)/\kappa_{2}(f)^{d}$ and $\kappa_{d}(f)$ is given by (2.5) for all positive integers $d$. Furthermore, $(-1)^{d}\kappa_{2d}^{*}(f)$ is a non-positive real number by Corollary 17.
5.1 CGF characteristic functions
Consider the set of standardized characteristic functions of random variables associated to cyclotomic generating functions. As mentioned above, it suffices to consider only basic CGFs, so consider the set of all standardized CGF characteristic functions on the complex plane,
@@ 5.1 CGF characteristic functions
Given $\phi_{\mathcal{X}^{*}}(z)\in\mathcal{C}_{\operatorname{CGF}}$, let $f(q)\in\Phi^{+}$ be the associated CGF, and let $\mathcal{X}$ be the corresponding random variable with $\mathbb{E}[q^{\mathcal{X}}]=f(q)/f(1)$, mean $\mu$ and standard deviation $\sigma$. Let $\tilde{f}(q)=q^{-\mu}f(q)/f(1)\in\mathbb{R}_{\geqslant 0}[q^{\pm 1/2}]$. Then, by (5.2)
$\phi_{\mathcal{X}^{*}}(z)=e^{-iz\mu/\sigma}f(e^{iz/\sigma})/f(1)=\tilde{f}(e^{iz/\sigma}).$ (5.4)
@@
From Montel’s theorem in complex analysis, $\mathcal{C}_{\operatorname{CGF}}$ is a normal family of entire functions if and only if it is bounded on all complex disks $|z|\leqslant R$. Hence we will bound $|\phi_{\mathcal{X}^{*}}(z)|$ in terms of $R$. Note that $f(q)=1$ if and only if $\sigma=0$ by Lemma 30(iii), in which case $\phi_{\mathcal{X}^{*}}(z)=1$ is bounded by any function greater than 1. So, we will assume $f(q)$ is not constant and $\sigma>0$. By [HZ15, Lem. 2.8], for all real $t\geqslant 0$,aaa[HZ15, Lem. 2.8] is incorrectly stated for all $t\in\mathbb{R}$, though one of the last steps in the argument requires $t\geqslant 0$.
$\mathbb{E}[e^{t\mathcal{X}^{*}}]\leqslant\exp\left(\frac{3}{2}t^{2}e^{2t/\sigma}\right).$ (5.5)
Since $\mathbb{E}[e^{z\mathcal{X}^{*}}]=\phi_{\mathcal{X}^{*}}(-iz)$, we can use this inequality to bound $|\phi_{\mathcal{X}^{*}}(z)|$ for all complex $|z|\leqslant R$ as follows.
For all $|z|\leqslant R$, we claim that $|\tilde{f}(e^{iz/\sigma})|\leqslant 2\tilde{f}(e^{R/\sigma})$. Indeed, since $f(q)\in\Phi^{+}$, we have $\tilde{f}(q)=\sum_{k=-N}^{N}a_{k}q^{k/2}$ for $a_{k}\geqslant 0$ satisfying $a_{k}=a_{-k}$. Therefore,
$\left|\tilde{f}(e^{iz/\sigma})\right|$ $=\left|\sum_{k=-N}^{N}a_{k}e^{izk/2\sigma}\right|\leqslant\sum_{k}a_{k}\left|e^{izk/2\sigma}\right|=\sum_{k}a_{k}e^{\operatorname{Re}(izk/2\sigma)}$
$\leqslant\sum_{k}a_{k}e^{R|k|/2\sigma}\leqslant\sum_{k}a_{k}(e^{Rk/2\sigma}+e^{-Rk/2\sigma})=\sum_{k}a_{k}e^{Rk/2\sigma}+\sum_{k}a_{-k}e^{-Rk/2\sigma}$
@@
By Lemma 30(iii), $\sigma^{2}\geqslant 1/4$ is bounded away from 00 since we assumed $f(q)$ is not constant. Therefore, by the claim, (5.4), and (5.5), we have the required uniform bound
$|\phi_{\mathcal{X}^{*}}(z)|=|\tilde{f}(e^{iz/\sigma})|\leqslant 2\tilde{f}(e^{R/\sigma})=2\phi_{\mathcal{X}^{*}}(-iR)=2\mathbb{E}[e^{R\mathcal{X}^{*}}]\leqslant 2\exp\left(\frac{3}{2}R^{2}e^{8R}\right).$
Theorem 40 (Converse of Frechét–Shohat for CGFs).
Suppose $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ is a sequence of random variables corresponding to cyclotomic generating functions such that $\mathcal{X}_{n}^{*}\Rightarrow\mathcal{X}$ for some random variable $\mathcal{X}$. Then $\mathcal{X}$ is determined by its moments and, for all $d\in\mathbb{Z}_{\geqslant 1}$,
@@
By Theorem 39, we may pass to a subsequence and assume $\phi_{\mathcal{X}_{n}^{*}}(z)$ converges uniformly on compact subsets of $\mathbb{C}$. Hence, they converge to an entire function $\phi(z)$. On the other hand, by Lévy continuity, $\phi_{\mathcal{X}_{n}^{*}}(t)\to\phi_{\mathcal{X}}(t)$ pointwise for all $t\in\mathbb{R}$. Thus, $\phi_{\mathcal{X}}(t)=\phi(t)$ for $t\in\mathbb{R}$.
@@
We may now prove Theorem 8 from the introduction.
Proof of Theorem 8.
Suppose $\mathcal{X}_{N}^{*}\Rightarrow\mathcal{X}$ converges in distribution. Then the result follows by Theorem 40. Conversely, suppose $\lim_{N\to\infty}\mu_{d}^{\mathcal{X}_{N}^{*}}$ exists and is finite for all $d\in\mathbb{Z}_{\geqslant 1}$. By Theorem 39, we may pass to a subsequence on which $\phi_{\mathcal{X}_{N}^{*}}(z)$ converges uniformly on compact subsets of $\mathbb{C}$. The limiting entire function $\phi(z)$ then has power series coefficients determined by the $\mu_{d}:=\lim_{N\to\infty}\mu_{d}^{\mathcal{X}_{N}^{*}}$, so the limit is independent of the subsequence we passed to and $\lim_{N\to\infty}\phi_{\mathcal{X}_{N}^{*}}(t)=\phi(t)$ for all $t\in\mathbb{R}$. Moreover, $\phi(t)$ is continuous at 00, so there exists $\mathcal{X}$ such that $\mathcal{X}_{N}^{*}\Rightarrow\mathcal{X}$ by [Bil95, Cor. 26.1]. ∎
5.2 Formal cumulants and cumulant generating functions
Recall formal cumulants from Section 2.3. The formal cumulants of cyclotomic polynomials have the following explicit growth rate.
Lemma 42.
For each fixed $d\geqslant 1$, the formal cumulants of the cyclotomic polynomials for $n>1$ satisfy
$\frac{(2d)!}{d}\left(\frac{n}{2\pi}\right)^{2d}\leqslant|\kappa_{2d}|\leqslant\frac{|B_{2d}|n^{2d}}{2d}.$ (5.6)
@@ 5.2 Formal cumulants and cumulant generating functions
See [Wik23] for a sketch of the proof. Applying the expression for the zeta function at positive even integers gives
@@
$\geqslant\frac{|B_{2d}|}{2d}\frac{1}{\zeta(2d)}$
@@
While cyclotomic generating functions are typically given in rational form, Lemma 16 allows their cumulants to be described in terms of the cyclotomic form as follows. One may for instance use Example 26 to describe the $q$-hook formula in cyclotomic form combinatorially.
Corollary 43.
If $f(q)=\prod_{j=1}^{k}\Phi_{n_{j}}(q)$, then for each $d\geqslant 1$ we have
@@
The asymptotic behavior of sequences of CGF random variables can be determined by their standardized cumulant generating function. For instance, $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ is asymptotically normal if and only if the corresponding sequence of standardized cumulants $(\kappa_{d}^{(N)})^{*}\to 0$ for all $d\geqslant 3$ as $N\to\infty,$ in which case by (5.3), $\log\phi_{\mathcal{X}_{N}^{*}}(t)\to-t^{2}/2$ for all $t\in\mathbb{R}$ as $N\to\infty$. Standardized second characteristic functions of CGFs are particularly well-behaved in the following sense.
Proposition 44.
@@
$\log\phi_{\mathcal{X}^{*}}(t)\leqslant-\frac{t^{2}}{2}=\log\phi_{\mathcal{N}(0,1)}(t).$ (5.7)
@@
By Theorem 39, $\phi_{\mathcal{X}}(z)$ is entire, so the singularities of $\log\phi_{\mathcal{X}}(z)$ come about from the zeros of $\phi_{\mathcal{X}}(z)$. To determine the zeros, it suffices to assume $\mathcal{X}$ is determined by $f(q)=\frac{\prod_{k=1}^{m}[a_{k}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}\in\Phi^{+}$ since the results pertain to the standardized random variable corresponding to $\mathcal{X}$. Furthermore, since $\phi_{\mathcal{X}}(z)=1$ if $f(q)=1$, which satisfies both claims, we may assume that $m=\max\{a_{k}\}>1$ and $m$ does not appear in the denominator multiset. Hence, $\Phi_{m}(q)$ is a factor of $f(q)$. Furthermore, from the product formula (2.14), we observe that the closest zero of $\phi_{\mathcal{X}}(z)=f(e^{iz})$ to the origin occurs at $z=2\pi/m$. Hence $\phi_{\mathcal{X}}(z)$ is analytic on the simply connected domain $|z|<2\pi/m$, so $\log\phi_{\mathcal{X}}(z)$ is analytic on the disk $|z|<2\pi/m$.
@@
Since $\Phi_{m}(q)$ is a factor of $f(q)$, we have $0<\kappa_{2}(\Phi_{m})\leqslant\sigma^{2}$ by Lemma 42 and Corollary 43. By the lower bound in (5.6), $\sqrt{\kappa_{2}(\Phi_{m})}\geqslant\frac{m}{\pi\sqrt{2}}$. Hence if $|z|<\sqrt{2}$,
@@
$\log\phi_{\mathcal{X}^{*}}(z)=-\sum_{d\geqslant 1}\frac{(-1)^{d-1}\kappa_{2d}}{(2d)!}\left(\frac{z}{\sigma}\right)^{2d}$ (5.8)
to compute $\log\phi_{\mathcal{X}^{*}}(t)$ for real $-\sqrt{2}<t<\sqrt{2}$. The inequality (5.7) comes from truncating this power series expansion after the first term and applying Corollary 17 to see that all terms on the right side are nonpositive. ∎
6 CGF monoids and related open problems
As mentioned in Definition 3, the set of basic cyclotomic generating functions forms a monoid under multiplication, $\Phi^{+}$. We also consider the larger monoid generated by all cyclotomic polynomials. Since many families of polynomials of interest are also either unimodal or log-concave with no internal zeros, we consider these submonoids as well, together with another variant using the Gale order on multisets. We conclude with a monoid associated to Hilbert series of polynomial rings quotiented by regular sequences and several open problems.
6.1 Basic, unimodal, and log-concave CGF monoids
Definition 45.
@@ 6.1 Basic, unimodal, and log-concave CGF monoids
Lemma 46.
We have $\Phi^{\operatorname{lcc}}\subset\Phi^{\operatorname{uni}}\subset\Phi^{+}\subset\Phi^{\pm}$, and for each monoid $\mathcal{M}$ in $\{\Phi^{\operatorname{lcc}},\Phi^{\operatorname{uni}},$ $\Phi^{+},\Phi^{\pm}\}$ the following facts hold.
@@
For (iv), first note that $f(1)\geqslant 1\neq 0$ for $f\in\Phi^{+}$, so $\Phi_{1}(q)=q-1$ is not a factor of $f$. Further recall that $\Phi_{n}(0)=\Phi_{n}(1)=1$ for $n\geqslant 2$ not a prime power. Hence, if $f$ had no factors of the form $\Phi_{p^{k}}(q)$, we would have $f(0)=f(1)=1$, forcing $f(q)=1$ since the coefficients of $f$ are nonnegative. The forwards implication in (v) is given by (iv), and the backwards implication follows from the fact that
@@
Finally, for (vi), $\varphi(n)$ is even for all $n\geqslant 3$, so any product of cyclotomic polynomials of odd degree must contain $\Phi_{2}$. ∎
Example 47.
@@
Intuitively, multiplying by $\Phi_{p}(q)=1+q+\cdots+q^{p-1}$ for $p$ prime tends to smooth out chaotic coefficients. Computationally, it appears to be necessary to include such a factor in unimodal or log-concave CGFs. More precisely, we conjecture the following analogue of Lemma 46(iv), which has been checked up to degree $50$.
Conjecture 48.
@@
We use the cyclotomic form for basic CGFs and the fact that we know a lower bound for the degree of the cyclotomic polynomial $\Phi_{n}(q)$ to find all basic CGFs of a given degree. Some initial data and OEIS identifiers [OEI23] can be found in Section 7. The sequence corresponding to $\Phi_{n}^{\pm}$ [OEI23, A120963] was created in 2006. We added the corresponding pages for the other cyclotomic monoids, so we believe the other cyclotomic monoids have not been studied in the literature.
6.2 Gale order and the associated CGF monoid
Given two multisets of the same size, say $A=\{a_{1}\leqslant a_{2}\leqslant\cdots\leqslant a_{m}\}$ and $B=\{b_{1}\leqslant b_{2}\leqslant\cdots\leqslant b_{m}\}$ both sorted into increasing order, we say $A\preceq B$ in Gale order provided $a_{k}\leqslant b_{k}$ for all $1\leqslant k\leqslant m$. This partial order is known by many other names; we are following [ARW16] for consistency. Gale studied this partial order in the context of matroids on $m$-subsets of $\{1,2,\dots,n\}$ in the 1960’s [Gal68].
Definition 51.